OpenStudy (anonymous):

find the slope,if it exists, of the line containing the pair of points (8,0) and (0,-8)

OpenStudy (anonymous):

\[{y_2-y_1\over x_2-x_1}\]

OpenStudy (anonymous):

slope k=y/x y=(-8-0)=-8 x=(0-8)=-8 so k=-8/-8=1

OpenStudy (anonymous):

is that answer simplify

OpenStudy (anonymous):

so m=1

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

same question but (1,-7) and (-1,-7)

OpenStudy (anonymous):

y=-7-(-7)=0 x=-1-1=-2 so k=0/-2=0

OpenStudy (anonymous):

so does that mean the slope is undefined